OpenStudy (anonymous):

Will medal. What is the common ratio, r , for the geometric sequence below? A1=-1, and An=4/3An=1 for n=2,3,4 ...

OpenStudy (anonymous):

\(A_n = \dfrac{4}{3}A_{n-1}\) , right?

OpenStudy (anonymous):

yes @OOOPS

OpenStudy (anonymous):

hehehe... only number \(\dfrac{4}{3}\) on it, so, it is ratio. done.

OpenStudy (anonymous):

Dat sit!! you asked :"What is the common ratio", and I answered that is \(\dfrac{4}{3}\), Is it not enough??